OpenStudy (anonymous):

Please help me real quick. :) Write the equation of the line that passes through the points (–1, –4) and (2, 5) in point-slope form.

OpenStudy (anonymous):

@wio @kohai

OpenStudy (luigi0210):

Use the slope formula and then the point-slope formula.

OpenStudy (nincompoop):

don't do it @babikirabdelmonim

OpenStudy (anonymous):

y-y1=m(x-x1)

OpenStudy (anonymous):

m = y2-y1/(x2-x1)

OpenStudy (anonymous):

where m is your slope

OpenStudy (anonymous):

slope= rise = y2-y1 = 5- (-4) = 5+4 = 9 = 3 = 3 ____ ______ _______ _______ ________ _____ run x2 - x1 2- (-1) 2+1 3 1

OpenStudy (anonymous):

i SWARE THAT'S THE CORRECT ANSWER, BY THE WAY =3

OpenStudy (anonymous):

swear

OpenStudy (anonymous):

to me

OpenStudy (anonymous):

I SWEAR

OpenStudy (anonymous):

SOME OF US HERE ARE GENIUS, DAAAH

OpenStudy (nincompoop):

na na na na na na nananaa .... na na

OpenStudy (anonymous):

http://www.youtube.com/watch?v=tZD4QsIJqiI 0:36

OpenStudy (nincompoop):

http://youtu.be/tZD4QsIJqiI?t=36s

OpenStudy (nincompoop):

na na na nananana nana.... na na

OpenStudy (kc_kennylau):

Find the slope of the line first, using the formula \(\Large m=\frac{y_2-y_1}{x_2-x_1}\)